# Green Finance and Carbon Emissions: A Nonlinear and Interaction Analysis Using Bayesian Additive Regression Trees

### v1

@@ Abstract
As a core policy tool for China in addressing climate risks, green finance plays a strategically important role in shaping carbon mitigation outcomes. This study investigates the nonlinear and interaction effects of green finance on carbon emission intensity (CEI) using Chinese provincial panel data from 2000 to 2022. The Climate Physical Risk Index (CPRI) is incorporated into the analytical framework to assess its potential role in shaping carbon outcomes. We employ Bayesian Additive Regression Trees (BART) to capture complex nonlinear relationships and interaction pathways, and use SHapley Additive exPlanations values to enhance model interpretability. Results show that the Green Finance Index (GFI) has a statistically significant inverted U-shaped effect on CEI, with notable regional heterogeneity. Contrary to expectations, CPRI does not show a significant impact on carbon emissions. Further analysis reveals that in high energy consumption scenarios, stronger green finance development contributes to lower CEI. These findings highlight the potential of green finance as an effective instrument for carbon intensity reduction, especially in energy-intensive contexts, and underscore the importance of accounting for nonlinear effects and regional disparities when designing and implementing green financial policies.
Keywords: green finance, carbon emissions, climate risk, BART, machine learning
@@ 1 Introduction
With the escalation of global climate change and the increasing frequency of extreme weather events, carbon reduction has emerged as an urgent global priority. Green finance, as a vital strategic instrument in addressing climate change, seeks to channel financial resources into environmentally sustainable projects aimed at improving energy efficiency and reducing carbon emissions wang2023can ; wu2024impact . It encompasses a range of financial instruments, including green credit, green bonds, and sustainable investment funds, which effectively allocate capital to projects that mitigate the adverse impacts of climate change niyazbekova2021growth . As the world’s largest carbon emitter, China has positioned green finance as a key driver in its transition toward a low-carbon economy, particularly in pursuit of its dual carbon targets: achieving carbon peaking by 2030 and carbon neutrality by 2060 lin2023china ; du2025decarbonization . According to the report Fintech Facilitates Green Finance in China: Cases and Outlook Institute2023Fintech , China has become the world’s largest green credit market and the second-largest green bond market, with the scale of green finance continuing to grow. However, the actual effectiveness of green finance is strongly shaped by local economic structures and environmental conditions. Recent studies suggest that climate risks significantly increase market uncertainty lee2024impact and influence investor decision-making behavior lu2025climate , thereby weakening the carbon reduction effects of green finance. Nevertheless, climate risk factors are rarely incorporated into analytical frameworks assessing the impact of green finance on carbon emissions. Moreover, most existing studies rely on traditional linear econometric models with manually specified interaction terms, and few have systematically examined the potential nonlinear effects or interaction mechanisms of green finance on carbon emissions.
@@
To address these research gaps, this study employs Bayesian Additive Regression Trees (BART), which are particularly suited to capturing nonlinear relationships and complex variable interactions. Drawing on panel data covering multiple economic and environmental indicators across 30 Chinese provinces from 2000 to 2022, the analysis incorporates climate risk into the modeling framework to systematically examine how green finance affects carbon emissions and whether climate risk has a significant impact on emission outcomes.
@@
This research makes three key contributions to bridge the gaps in variable integration and methodological innovation: first, it incorporates climate risk into the analytical framework to assess its direct effect on carbon emissions and its potential interaction effects with green finance; second, by employing BART, it overcomes the limitations of traditional linear econometric approaches in capturing nonlinear dynamics and interaction effects. Combined with interpretive tools such as Partial Dependence Plots (PDPs) and SHapley Additive exPlanations (SHAP), the study reveals the underlying mechanisms through which green finance affects carbon emissions and provides interpretable empirical evidence; third, methodologically, this study expands the analytical toolkit at the intersection of green finance and carbon emissions research, demonstrating the feasibility and value of machine learning approaches in environmental policy evaluation. It offers technical guidance for future research and provides a robust empirical foundation for the formulation of green transition policies.
@@ 2.1 Impact factors of carbon emissions
Extensive and mature research has been conducted to identify the driving forces of carbon emissions. Among the various analytical frameworks, the STIRPAT model york2003stirpat remains one of the most widely applied. Yang et al. yang2018what found a temporal shift in key emission drivers from urbanization to foreign trade. Tang et al. tang2024study further highlighted that in the metal smelting industry, population, coal consumption, and urbanization were the most sensitive factors. Xiao and Peng xiao2023decomposition , through Granger causality tests, confirmed that economic growth, urbanization, industrial structure, and energy structure significantly influence carbon emissions. To better address spatial and temporal variations, Chen et al. chen2021analysis adopted the GTWR-STIRPAT model and found that energy efficiency, trade, industrial structure, and environmental regulation exerted spatially heterogeneous and time-lagged effects. Similarly, Li & Wang li2022dynamic , using the Spatial Durbin Model and Panel Threshold Model, demonstrated that energy consumption and non-green technologies drive short-term increases in carbon emissions, whereas green innovation and industrial upgrading play a key role in long-term emission reductions. In addition, a number of researchers have employed decomposition approaches to examine the structural components of carbon emissions. Chen et al. chen2019driving identified several key driving factors, including potential carbon factor effects, energy intensity effects, economic scale effects, emission efficiency effects, and energy use efficiency effects. Liu et al. liu2019driving found that economic activity, emission efficiency, and potential carbon factors were the main contributors to emission growth, while improvements in energy intensity, pure technical efficiency, and scale efficiency helped mitigate emissions. Zhang et al. zhang2022carbon concluded that economic efficiency is the sole driver of industrial carbon emissions, while energy intensity remains the most critical factor in emission reduction. In contrast, Sun et al. sun2022characteristics found that energy structure and industrial structure have relatively minor effects on the growth rate of carbon emissions, and that energy intensity is the primary factor contributing to decoupling.
@@ 2.2 Green finance and carbon emissions
Green finance, as an important financial tool for promoting low-carbon transition, has been widely confirmed by empirical studies to have a significant suppressive effect on carbon emissions. Lin et al. lin2023impact found that green finance significantly reduces local carbon emissions and generates spillover effects in surrounding regions. Among various green financial instruments, Wang et al. wang2023can argue that the emission reduction effect of green credit is more pronounced than that of green bonds. Green finance not only contributes directly through investment in low-carbon projects but also facilitates carbon reduction indirectly by stimulating green technological innovation, optimizing industrial structure, and improving energy efficiency wang2022how ; tu2025impact ; ran2023driving .
However, the emission reduction effects of green finance are not always consistent across different contexts. To explore its underlying mechanisms more deeply, recent studies have incorporated contextual variables into their analyses. Zhao and Li zhao2024role found that in regions with high political risk or high economic complexity, the effectiveness of green finance in reducing emissions is significantly weakened, suggesting that the institutional environment plays a moderating role. Ma and Fei ma2024research further emphasized that factors such as the level of governmental environmental attention and the development of the financial sector exhibit marked heterogeneity in shaping the effectiveness of green finance in influencing carbon emissions.
Moreover, researchers have begun to explore the nonlinear nature of green finance’s impact. Sun sun2022characteristics and Yu et al. yu2024machine found that the effect of green finance on carbon emissions is not a simple linear decline. Instead, it varies significantly under different conditions of energy consumption structure and financial technology development. Similarly, Cai et al. cai2025how uncovered complex nonlinear interactions between green finance and other factors such as environmental investment, inclusive finance, and financial regulation, all of which significantly influence carbon emission intensity.
In summary, while the role of green finance in carbon emission reduction is increasingly evident, its underlying mechanisms exhibit notable multidimensionality, regional heterogeneity, and nonlinearity.
@@ 2.3 Application of machine learning algorithms in carbon emission
Existing literature has made significant progress in exploring the factors influencing carbon emissions using traditional econometric models, particularly the STIRPAT framework li2017industrial ; zhang2019identifying ; thio2022estimation . However, these approaches generally rely on assumptions of linear relationships among variables and are often constrained by researchers’ prior selection of explanatory factors. As a result, they have limitations in capturing complex nonlinear relationships and high-order interaction effects. In recent years, the development of machine learning methods has provided a novel technical pathway for modeling carbon emission mechanisms and identifying key drivers. Compared to traditional models, machine learning algorithms can automatically learn nonlinear mapping relationships from large-scale datasets and exhibit greater robustness and adaptability kim2022novel ; vazquez2022automatic . As machine learning continues to expand across various domains, its application in carbon emission analysis has also gained increasing traction. For instance, various studies have applied machine learning models, such as Random Forest, Decision Trees, and XGBoost qin2022estimation ; shan2025unveiling , to identify the key drivers of carbon emissions from large sets of candidate variables, often ranging from dozens to hundreds. These approaches underscore the advantages of machine learning over traditional linear models, particularly in managing high-dimensional, complex systems and capturing nonlinear relationships and interaction effects among variables qin2022estimation ; ahmed2022influencing ; gao2023approach . To address the “black-box" nature of these machine learning algorithms, some studies have further introduced approaches to improve the interpretability of results, such as SHAP (Shapley Additive Explanations) and ALE (Accumulated Local Effects). Gao et al. gao2023approach and Shan et al. shan2025unveiling demonstrate how these methods can reveal the marginal contributions of variables and identify nonlinear “tipping points”, thereby enhancing the interpretability of model outputs and their value for policy guidance.
@@ 3 Data
This research constructs a panel dataset that covers 30 Chinese provinces over the period 2000–2022. All variables are measured at the provincial level, with each variable observed separately for each province, allowing the investigation of how provincial characteristics influence carbon emission outcomes. A more detailed description of the 136 variables is provided in the following subsections.
3.1 Variable description
3.1.1 Dependent variable
We employ carbon emission intensity (CEI) as the core indicator to measure regional carbon emission performance. CEI is defined as the amount of carbon emissions per unit of GDP, reflecting the carbon efficiency of regional economic activities.
Compared to total carbon emissions, CEI controls for differences in regional economic scale, providing a more accurate measure of green development performance across regions. This makes it more suitable for analyzing the relationship between green finance development and regional carbon emissions, as it better reflects the environmental efficiency of economic activities under varying economic contexts.
Following the method proposed by Wang et al. wang2022digital , this study adopts the natural logarithm of carbon emissions per unit of GDP as the response variable. The specific formula is as follows:
@@ 3.1.1 Dependent variable
3.1.2 Core independent variables
Referring to Lee and Lee LEE2022105863 , Ran et al. ran2023driving and Wu et al. wu2024impact , this study constructs a composite green finance index to measure the level of green finance development across regions. The index is based on seven secondary indicators: green bond, green fund, green credit, green insurance, green equity, green support, and green investment. These sub-indicators are first standardized to ensure cross-provincial comparability, and then aggregated using the entropy weight method to generate the final composite index. A higher index value indicates a higher level of green finance development in the corresponding region. Table 1 provides additional information regarding these secondary indicators.
@@ 3.1.2 Core independent variables
Primary indicators Secondary indicators Variable declaration
Green Financial Index (GFI) Green bond Total green bond issuance / all bond issuance
Green fund Market Value of Green Funds / Total Market Value of All Funds
Green credit Credit for Environmental Protection Projects / Total Provincial Credit
@@
Green insurance Revenue from Environmental Liability Insurance / Total Insurance Premium Income
@@
Green equity Trading Volume of Carbon Emissions, Energy Use Rights, and Pollution Discharge Rights / Total Volume of Equity Market Transactions
@@
Green support Fiscal Expenditure on Environmental Protection / General Public Budget Expenditure
Green investment Environmental pollution control investment / GDP
3.1.3 Other independent variables
Existing literature on the drivers of carbon emissions from an energy perspective has primarily focused on dimensions such as energy structure, energy efficiency, and energy intensity chen2019driving ; liu2019driving ; zhang2022carbon ; xiong2025emission . Meanwhile, total energy consumption (TEC) has been widely recognized as a key factor contributing to the increase in carbon emissions and has become a general consensus in the academic community li2022dynamic ; yu2024machine . However, when shifting the focus to carbon emission intensity, the impact of TEC becomes more complex. zhao2024role . Therefore, this paper includes TEC as one of the explanatory variables. On the one hand, it aims to further examine the impact of TEC on CEI. On the other hand, it seeks to explore whether green finance interacts with TEC in a way that facilitates the transformation of the energy consumption structure and ultimately contributes to the reduction of carbon emission intensity.
@@ 3.1.3 Other independent variables
Climate change has become a central concern globally, and managing climate risk exposure through investment strategies has emerged as a key decision-making factor for investors lu2025climate . Climate risk, particularly physical climate risk, can increase market uncertainty and reduce investment incentives, thereby inhibiting the development of green finance markets. In this context, drawing on the methodology proposed by Guo et al. guo2024impact , this study constructs a Climate Physical Risk Index (CPRI) based on region-specific meteorological observations and incorporates it as one of the explanatory variables in the model. To capture both the intensity and frequency of extreme weather events, the CPRI evaluates physical climate risk across four dimensions: extreme low temperature, extreme high temperature, extreme rainfall, and extreme drought. Based on daily meteorological data from local stations, the annual number of days exceeding each threshold is calculated for every province. To ensure comparability across provinces, each of the four sub-indices is standardized. The overall CPRI is then computed as the weighted average of these standardized sub-indices. A higher CPRI value indicates greater exposure to physical climate risk in a given region.

### v2

@@ Abstract
This study investigates the nonlinear and interaction effects of green finance on carbon emission intensity (CEI) using Chinese provincial panel data from 2000 to 2022. The Climate Physical Risk Index (CPRI) is incorporated into the analytical framework to assess the potential influence of climate risk on carbon emission performance. We employ Bayesian Additive Regression Trees (BART) to capture complex nonlinear relationships and interaction pathways between the variables, then use SHapley Additive exPlanations to enhance model interpretability. Results show that the Green Finance Index (GFI) has a significant effect on CEI, exhibiting an inverted U-shaped effect, with notable regional heterogeneity. Contrary to expectations, there is insufficient evidence to suggest a significant association between CPRI and CEI. Further analysis reveals that in high energy consumption scenarios, stronger green finance development contributes to lower CEI. These findings highlight the potential of green finance as an effective instrument for carbon reduction, especially in energy-intensive contexts, and underscore the importance of accounting for nonlinear effects and regional disparities when designing and implementing green financial policies.
Keywords: green finance, carbon emissions, climate risk, Bayesian Additive Regression Trees
@@ 1 Introduction
With the escalation of global climate change and the increasing frequency of extreme weather events, reducing carbon emissions has emerged as an urgent global priority. Green finance, as a vital strategic instrument in addressing climate change, seeks to channel financial resources into environmentally sustainable projects aimed at improving energy efficiency and reducing carbon emissions wang2023can ; wu2024impact . Green finance encompasses a range of financial instruments, including green credit, green bonds, and sustainable investment funds, which effectively allocate capital to projects that mitigate the adverse impacts of climate change niyazbekova2021growth . As the world’s largest carbon emitter, China has positioned green finance as a key driver in its transition toward a low-carbon economy, particularly in pursuit of its dual carbon targets: achieving carbon peaking by 2030 and carbon neutrality by 2060 lin2023china ; du2025decarbonization . According to the report Institute2023Fintech , China has become the world’s largest green credit market and the second-largest green bond market, with the scale of green finance continuing to grow. However, the actual effectiveness of green finance is strongly shaped by local economic structures and environmental conditions. Recent studies suggest that climate risks significantly increase market uncertainty lee2024impact and influence investor decision-making lu2025climate , thereby affecting the carbon reduction effects of green finance. Nevertheless, climate risk factors are rarely incorporated into analytical frameworks assessing the impact of green finance on carbon emissions. Moreover, most existing studies rely on traditional linear econometric models with manually specified interaction terms, and few have systematically examined the potential nonlinear effects or interaction mechanisms of green finance on carbon emissions.
@@
To address these research gaps, this study employs Bayesian Additive Regression Trees (BART), which is particularly suited to capture nonlinear relationships and complex variable interactions. Drawing on panel data covering multiple economic and environmental indicators across 30 Chinese provinces from 2000 to 2022, the analysis incorporates climate risk into the modeling framework to systematically examine how green finance affects carbon emissions and whether climate risk has a significant impact on emission outcomes.
@@
This research makes three key contributions to bridge the gaps in variable integration and methodological innovation: first, it incorporates climate risk into the analytical framework to assess its direct effect on carbon emissions and its potential interaction effects with green finance; second, by employing BART, it overcomes the limitations of traditional linear econometric approaches in capturing nonlinear dynamics and interaction effects. Leveraging interpretable tools such as Partial Dependence Plots (PDPs) and SHapley Additive exPlanations (SHAP), this study reveals the underlying mechanisms through which green finance associates with carbon emissions and provides interpretable model-based evidence; third, methodologically, this study expands the analytical toolkit at the intersection of green finance and carbon emissions research, demonstrating the feasibility and value of machine learning approaches in environmental policy evaluation. As a consequence, this contribution offers technical guidance for future research and provides a robust empirical foundation for the formulation of green transition policies.
@@ 2.1 Impact factors of carbon emissions
Extensive research has been conducted to identify the driving forces of carbon emissions. Among the various analytical frameworks, the STIRPAT model of york2003stirpat remains one of the most widely applied. yang2018what found a temporal shift in key emission drivers from urbanization to foreign trade. tang2024study further highlighted that in the metal smelting industry, population, coal consumption, and urbanization were the most sensitive factors. Through Granger causality tests, xiao2023decomposition confirmed that economic growth, urbanization, industrial structure, and energy structure significantly influence carbon emissions. To better address spatial and temporal variations, chen2021analysis adopted the GTWR-STIRPAT model and found that energy efficiency, trade, industrial structure, and environmental regulation exerted spatially heterogeneous and time-lagged effects. Similarly, li2022dynamic , using the Spatial Durbin Model and Panel Threshold Model, demonstrated that energy consumption and non-green technologies drive short-term increases in carbon emissions, whereas green innovation and industrial upgrading play a key role in long-term emission reductions. In addition, several researchers have employed decomposition analysis to examine the structural components of carbon emissions. For examples, chen2019driving identified several key driving factors, including potential carbon factor effects, energy intensity. liu2019driving found that economic activity, emission efficiency, and potential carbon factors were strongly associated with emissions growth, while improvements in energy intensity, pure technical efficiency, and scale efficiency helped mitigate emissions. zhang2022carbon concluded that economic efficiency is the sole driver of industrial carbon emissions, while energy intensity remains the most critical factor in carbon reduction. In contrast, sun2022characteristics found that energy and industrial structures have relatively minor effects on the growth rate of carbon emissions, and that energy intensity is the primary factor contributing to decoupling.
@@ 2.2 Green finance and carbon emissions
Green finance, as an important financial tool to promote the low-carbon transition, has been widely confirmed by empirical studies to have a significant suppressive effect on carbon emissions. lin2023impact found that green finance significantly reduces local carbon emissions and generates spillover effects in surrounding regions. Among various green financial instruments, wang2023can argue that the carbon reduction effect of green credit is more pronounced than that of green bonds. Green finance not only contributes directly through investment in low-carbon projects, but also facilitates carbon reduction indirectly by stimulating green technological innovation, optimizing industrial structure, and improving energy efficiency wang2022how ; tu2025impact ; ran2023driving .
However, the carbon reduction effects of green finance are not always consistent across different contexts. To explore its underlying mechanisms more deeply, recent studies have incorporated contextual variables into their analyses. zhao2024role found that in regions with high political risk or high economic complexity, the effectiveness of green finance in reducing emissions is significantly weakened, suggesting that the institutional environment plays a moderating role. ma2024research further emphasized that factors such as the level of governmental environmental attention and the development of the financial sector exhibit marked heterogeneity in shaping the effectiveness of green finance in influencing carbon emissions.
Moreover, researchers have begun to explore the nonlinear nature of green finance’s impact. sun2022characteristics and yu2024machine found that the effect of green finance on carbon emissions is not a simple linear decline. Instead, it varies significantly under different conditions of energy consumption structure and financial technology development. Similarly, cai2025how uncovered complex nonlinear interactions between green finance and other factors such as environmental investment, inclusive finance, and financial regulation, all of which significantly influence carbon emission intensity.
In summary, while the role of green finance in carbon emission reduction is increasingly evident, its underlying mechanisms exhibit notable complexity, regional heterogeneity, and nonlinearity.
@@ 2.3 Application of machine learning algorithms in carbon emission
The existing literature has made significant progress in exploring the factors influencing carbon emissions using traditional econometric models, particularly the STIRPAT framework li2017industrial ; zhang2019identifying ; thio2022estimation . However, these approaches generally rely on assumptions of linear relationships among variables and are often affected by researchers’ prior selection of explanatory factors. As a result, they have limitations in capturing complex nonlinear relationships and high-order interaction effects. In recent years, the development of machine learning methods has provided a novel technical pathway for modeling carbon emission mechanisms and identifying key drivers. Compared to traditional models, machine learning algorithms can automatically learn nonlinear mapping relationships from large-scale datasets and exhibit greater robustness and adaptability kim2022novel ; vazquez2022automatic . As machine learning continues to expand across various domains, its application in carbon emission analysis has also gained increasing traction. For instance, various studies have applied machine learning models, such as random forest, decision trees, and XGBoost qin2022estimation ; shan2025unveiling , to identify the key drivers of carbon emissions from large sets of candidate variables, often ranging from dozens to hundreds. These approaches underscore the advantages of machine learning over traditional linear models, particularly in managing high-dimensional, complex systems and capturing nonlinear relationships and interaction effects among variables qin2022estimation ; ahmed2022influencing ; gao2023approach . To address the “black-box" nature of these machine learning algorithms, some studies have further introduced approaches to improve the interpretability of results, such as SHAP (Shapley Additive Explanations) and ALE (Accumulated Local Effects). gao2023approach and shan2025unveiling , respectively, demonstrate how these methods can reveal the marginal contributions of variables and identify nonlinear “tipping points”, thereby enhancing the interpretability of model outputs and their value for policy guidance.
@@ 3 Data
This research constructs a panel dataset covering 30 Chinese provinces over the period 2000–2022. All variables are measured at the provincial level, which allows us to examine how provincial characteristics influence carbon emission outcomes. Due to data limitations, Tibet, Hong Kong, Macau, and Taiwan are excluded. The following subsections provide details on the variables, their sources, and preprocessing methods.
3.1 Dependent variable
This study employs carbon emission intensity (CEI) as the core indicator to measure regional carbon emission performance. CEI is defined as the amount of carbon emissions per unit of Gross Domestic Product (GDP), reflecting the carbon efficiency of regional economic activities.
Compared to total carbon emissions, CEI controls differences in regional economic scale, providing a more accurate measure of green development performance across regions. This makes it more suitable for analyzing the relationship between green finance development and regional carbon emissions, as it better reflects the environmental efficiency of economic activities under varying economic contexts.
Following the method proposed by wang2022digital , we adopt the natural logarithm of carbon emissions per unit of GDP as the response variable. The specific formula is as follows:
@@ 3.1 Dependent variable
3.2 Core independent variable
Referring to LEE2022105863 , ran2023driving and wu2024impact , we construct a composite green finance index as the core independent variable to measure the level of green finance development across regions. The index is based on seven secondary indicators: green bond, green fund, green credit, green insurance, green equity, green support, and green investment. These sub-indicators are first standardized to ensure comparability across provinces and then aggregated using the entropy value method of ding2016 to generate the final composite index. The entropy value method is employed to determine the weights of the sub-indicators objectively. In computing the weights for the sub-indicators, the sign (positive/negative) of each indicator signal is defined following wu2024impact ; the detailed calculation procedure is provided in LEE2022105863 . The entropy value method reduces potential bias and ensures that the resulting green finance index more accurately reflects the underlying data characteristics. A higher index value indicates a higher level of green finance development in the corresponding region. Table 1 provides additional information regarding these secondary indicators.
@@ 3.2 Core independent variable
Primary indicators Secondary indicators Weight Variable declaration
Green Financial Index (GFI) Green bond 0.147 Total Green Bond Issuance / All Bond Issuance
Green fund 0.144 Market Value of Green Funds / Total Market Value of All Funds
Green credit 0.130 Credit for Environmental Protection Projects / Total Provincial Credit
@@
Green insurance 0.129 Revenue from Environmental Liability Insurance / Total Insurance Premium Income
@@
Green equity 0.148 Trading Volume of Carbon Emissions, Energy Use Rights, and Pollution Discharge Rights / Total Volume of Equity Market Transactions
@@
Green support 0.162 Fiscal Expenditure on Environmental Protection / General Public Budget Expenditure
Green investment 0.140 Environmental Pollution Control Investment / GDP
3.3 Other independent variables
3.3.1 Total energy consumption
Total Energy Consumption (TEC) is an important factor affecting CEI. Higher TEC generally increases total carbon emissions (the numerator of CEI), while greater energy use can also stimulate economic activity and expand GDP (the denominator), which may partially offset its impact on CEI. Including TEC as an explanatory variable allows us to examine its overall effect on CEI and explore its potential interaction with green finance in promoting a low-carbon transition.
3.3.2 Climate physical risk index
@@
A higher climate physical risk may increase regional economic vulnerability to extreme weather, potentially affecting energy use, industrial output, and investment patterns, which in turn can influence CEI. In this context, drawing on the methodology proposed by guo2024impact , this study constructs a Climate Physical Risk Index (CPRI) based on region-specific meteorological observations and incorporates it as one of the explanatory variables in the model. To capture both the intensity and frequency of extreme weather events, the CPRI evaluates physical climate risk across four dimensions: extreme low temperature, extreme high temperature, extreme rainfall, and extreme drought. Based on daily meteorological data from local stations, the annual number of days exceeding each threshold is calculated for every province. In order to maintain comparability across provinces, the four sub-indices are rescaled to the same interval and then aggregated into a single value indicator. As a result, a higher CPRI value indicates greater exposure to physical climate risk in a given region.
3.3.3 Socioeconomic and environmental factors